A family has five children. The probablilty of having a girl is 1/2. What is the probability of having no more than 3 boys?
Answer by stanbon(48568)

A family has five children. The probablilty of having a girl is 1/2. What is the probability of having no more than 3 boys?
---------
P(0<=X<=3)= binomcdf(5,.5,3)=0.81
This from using a TI-83
another way is 1-[P(X=4)+P(X=5)]
P(x=4)=5C4(1/2)^5
P(x=5)=1/32
Cheers,
Stan H.